# Overlap-Aware Feature Learning for Robust Unsupervised Domain Adaptation for 3D Semantic Segmentation

### v1

@@ III-A Preliminaries and Overview
In the context of unsupervised domain adaptation on 3D point cloud semantic segmentation (PCSS-UDA) under adversarial conditions of the target domain, the labeled source domain is denoted as and the unlabeled target domain is denoted as . To obtain a point-wise classification of inputs, the pre-trained model is used to map the point cloud to the inference distributions of classes , which can be further divided into the feature embedding module and the classification module . The target model parameters learned by UDA are denoted as . For robustness evaluation, the target point cloud may suffer from subtle perturbations: , where denotes the applied perturbation. Our approach aims to achieve robust PCSS-UDA under adversarial conditions, including adversarial attacks and challenging weather conditions. We follow the recent pseudo-labeling technique, CosMix [19], which utilizes a teacher-student framework. The teacher model generates high-confidence pseudo-labels for unlabeled target data and gradually distills knowledge from a pre-trained source model. Our work consists of three key components: a robustness evaluation model, an Invertible Attention Alignment Module with overlap suppression loss (IAAM) and a Quality-guided Contrastive Memory Bank (QC-MB).
@@ III-B Robustness Evaluation Model
Adversarial attacks. Adversarial attacks typically perturb the original inputs with subtle perturbations but significantly degrade the deep learning model with high confidence [24, 25], which can be divided into white-box attacks and black-box attacks. White-box attacks utilize the parameters of the trained model and data labels to generate adversarial perturbations. In contrast, black-box attacks rely on the transferability of adversarial examples while limited for different model architectures and training datasets [26, 27, 28]. To accurately assess the worst-case performance degradation under input perturbations, we incorporate the white-box attack with the ground truth target labels and the PCSS-UDA model to generate target adversarial examples while target labels are unavailable during training. For point cloud input , and its labels , the adversarial objective is to generate adversarial examples subject to . In this work, we adopt the projected gradient descent (PGD) attack [9] and the iterative fast gradient sign descent (I-FGSM) attack [22] as the adversarial objective function. The generation of I-FGSM adversarial examples is denoted as:
in contrast, PGD extends I-FGSM by introducing an additional projection step, ensuring that adversarial perturbations remain within a bounded norm constraint. This prevents perturbations from growing arbitrarily large over iterations. The update rule for PGD attack is denoted as:
in which is the adversarial loss function, typically using the cross-entropy loss for classification tasks.
Robustness metric. As illustrated in Fig. 2, different perturbation intensities cause varying degradation of model performance. To quantitatively evaluate the robustness of the PCSS-UDA model, the robustness score is denoted as:
where the Mean Intersection over Union (mIoU) is a widely used evaluation metric in semantic segmentation, which is computed as:
where is the number of classes, and represent the predicted and ground truth sets for class , respectively. The robustness drop metric quantifies the relative performance degradation under adversarial perturbations by comparing mIoU on clean and adversarial examples.
@@ III-C Invertible Attention Alignment Module (IAAM)
Flow-based bidirectional alignment. To effectively leverage the intrinsic feature structure of the target domain, we employ a flow-based bidirectional alignment mechanism. The INN model learns a bidirectional feature mapping between the source and the target domains using INN with forward () and inverse () transformations. This process not only preserves the structural integrity of the target domain, but also enhances feature representations by integrating mapped features with the original representations. The mapping is optimized as follows:
where denotes the mapping loss of features. Specifically, we utilize , which denotes the optimal transport-based Wasserstein distance [29, 30, 31] computed through Sinkhorn iterations [32]. To preserve semantic integrity during cyclic transformations, we introduce the cyclic consistency loss:
where represents the -norm (empirically set to in our experiments).
Attention-guided pseudo-label generation. Inspired by earlier HMA approaches [17] in image classification, we utilize multi-head attention module to generate domain-invariant pseudo-labels and class distributions for mapping features through:
where are query, key, and value matrices derived from domain-transformed features. The attention training combines:
where denotes the cross-entropy loss, and is the classifier output of student model for target domain features. The attention module ensures distribution consistency across original and transformed features. The complete optimization for INN training integrates the mentioned bidirectional alignment and attention generation losses:
this joint optimization enables simultaneous feature alignment and feature-aware pseudo-label generation.
Overlap suppression loss. To resolve feature overlap for different classes, we propose the overlap loss:
where
this improves class prediction accuracy for joint feature distributions, and
in which represents the prediction probability for the ground truth label of the -th sample and denotes the predicted probability for the other classes of the -th sample. This loss penalizes inter-class feature overlap. The temperature parameter controls overlap sensitivity, with higher values increasing inter-class margin enforcement. The hyper-parameter balances label matching precision and class overlap suppression.
In this work, we refine the pseudo-label training with original adaptation loss and overlap loss during the UDA training, which helps minimize feature overlap for different classes. As illustrated in Fig. 3, the attention-guided pseudo-labeling module is used to predict class distributions and pseudo-labels of the mapping features , which are complementary to source domain features. In contrast, the labels of mapping features remain identical to source features to guide overlap learning across domains.
@@ III-D Quality-Aware Contrastive Memory Bank
Quality-guided sampling. We utilize the quality-guided sampling module to dynamically filter out noise and prioritize high-quality features in the latent space. To compute the quality scores, we take into account both the density and cleanliness scores, which can be denoted as:
where retrieves k nearest neighbors in the latent space and balances density and cleanliness scoring which is empirically set to 1.0 in our experiment. Specifically, the density score quantifies the local concentration of similar features to promote robust clusters, while the cleanliness score penalizes the presence of neighboring features from different classes to ensure feature purity. For imbalance class weight input, we adaptively select neighbors by class-aware scaling:
where is the class population, the adaptively increases for head classes while reducing for tail classes, mitigating class imbalance in different scenarios.
Momentum contrastive learning. To further improve the quality of pseudo-label generation and improve the discriminability of class-wise features, we introduce a contrastive learning memory bank. This memory bank ensures that the model learns more structural and high-confidence representations by storing high-quality class-specific anchor features. Similar to [6], selected features undergo a projection module before memory integration. The memory bank maintains prototypes per class through momentum updating, with updating rate and class-wise updating embeddings :
The contrastive learning objective enforces structural feature discrimination through:
in which denotes the cosine similarity of the p-th positive prototype. In contrast, denotes the cosine similarity of the q-th negative prototype. The temperature parameter scales the logits before the softmax operation, controlling the sharpness of the similarity distribution. Empirically, we set to balance feature separation and stability in contrastive learning.
As shown in Fig. 4, the contrastive learning module is used to pull the embedding inputs of the same classes closer to the stored anchor embeddings while pushing away from others, encouraging learning from high-quality and more discriminative features. In the update stage, the memory bank is gradually updated with the high-quality fused features and , note that the student classifier is used to predict pseudo-labels for the mapping features and guide high-quality sampling during the update stage of the memory bank. In the contrastive learning stage, the mapping feature is guided to align with high-quality stored anchor features and combined with the supervised segmentation loss during adaptation. The overall objective function for the memory bank training is:
where the hyper-parameter balances the contrastive learning and semantic segmentation learning.
@@ IV-A Implementation Details
We use SynLiDAR as the synthetic dataset for synthetic-to-real adaptation and SemanticPOSS as the real-world dataset. SynLiDAR is a large-scale synthetic dataset comprising 198,396 annotated point clouds with 32 semantic classes, while SemanticPOSS consists of 2,988 annotated real-world point clouds and 14 semantic labels. For real-to-real domain adaptation, we employ SemanticKITTI, a widely used large-scale dataset that provides point-wise semantic annotations of the KITTI dataset [36]. In addition, we use the SemanticSTF dataset, which contains 2,076 manually annotated scans captured by the Velodyne HiDL64S3D LiDAR sensor from STF [37] dataset. SemanticSTF includes various adverse weather conditions, such as snow, dense fog, light fog, and rain. Following CosMix [19], we adopt the same training and validation split for , while for , we use the original validation split. We employ the MinkUnet34 3D sparse convolution model [2] as our backbone architecture. The hyper-parameter and are empirically set to 1.0 and 20.0 for overlap suppression. During contrastive learning, we empirically set the number of class anchors stored in the contrastive memory bank to 64 to optimize memory usage and computational efficiency and set to 0.98 for memory update. All experiments are conducted using the PyTorch [38] framework on 2 NVIDIA GeForce RTX4090.

### v2

@@ III-A Preliminaries and Overview
In the context of unsupervised domain adaptation on 3D point cloud semantic segmentation (PCSS-UDA) under adversarial conditions of the target domain, the labeled source domain is denoted as and the unlabeled target domain is denoted as . To obtain a point-wise classification of inputs, the pre-trained model is used to map the point cloud to the inference distributions of classes , which can be further divided into the feature embedding module and the classification module . The target model parameters learned by UDA are denoted as . For robustness evaluation, the target point cloud may suffer from subtle perturbations: , where denotes the applied perturbation.
Our approach aims to achieve robust PCSS-UDA under adversarial conditions, including adversarial attacks and challenging weather conditions. We follow the recent pseudo-labeling technique, CosMix [19], which utilizes a teacher-student framework. The teacher model generates high-confidence pseudo-labels for unlabeled target data and gradually distills knowledge from a pre-trained source model. Our work consists of three key components: a robustness evaluation model, an Invertible Attention Alignment Module with overlap suppression loss (IAAM) and a Quality-guided Contrastive Memory Bank (QC-MB).
@@ III-B Robustness Evaluation Model
Adversarial attacks. Adversarial attacks typically perturb the original inputs with subtle perturbations but significantly degrade the deep learning model with high confidence [24, 25], which can be divided into white-box attacks and black-box attacks. White-box attacks utilize the parameters of the trained model and data labels to generate adversarial perturbations. In contrast, black-box attacks rely on the transferability of adversarial examples while limited for different model architectures and training datasets [26, 27, 28]. To accurately assess the worst-case performance degradation under input perturbations, we incorporate the white-box attack with the ground truth target labels and the PCSS-UDA model to generate target adversarial examples while target labels are unavailable during training. For point cloud input , and its labels , the adversarial objective is to generate adversarial examples subject to . In this work, we adopt the projected gradient descent (PGD) attack [9] and the iterative fast gradient sign descent (I-FGSM) attack [22] as the adversarial objective function.
@@ III-C Invertible Attention Alignment Module (IAAM)
Flow-based bidirectional alignment. To effectively leverage the intrinsic feature structure of the target domain, we employ a flow-based bidirectional alignment mechanism. The INN model learns a bidirectional feature mapping between the source and the target domains using INN with forward () and inverse () transformations. This process not only preserves the structural integrity of the target domain, but also enhances feature representations by integrating mapped features with the original representations.
Attention-guided pseudo-label generation. Inspired by earlier HMA approaches [17] in image classification, we utilize multi-head attention module to generate domain-invariant pseudo-labels and class distributions for mapping features. The attention module ensures distribution consistency across original and transformed features. The complete optimization for INN training integrates the mentioned bidirectional alignment and attention generation losses, which enables simultaneous feature alignment and feature-aware pseudo-label generation.
Overlap suppression loss. To resolve feature overlap for different classes, we propose the overlap loss. This loss penalizes inter-class feature overlap. The temperature parameter controls overlap sensitivity, with higher values increasing inter-class margin enforcement. The hyper-parameter balances label matching precision and class overlap suppression.
In this work, we refine the pseudo-label training with original adaptation loss and overlap loss during the UDA training, which helps minimize feature overlap for different classes. The attention-guided pseudo-labeling module is used to predict class distributions and pseudo-labels of the mapping features , which are complementary to source domain features. In contrast, the labels of mapping features remain identical to source features to guide overlap learning across domains.
@@ III-D Quality-Aware Contrastive Memory Bank
Quality-guided sampling. We utilize the quality-guided sampling module to dynamically filter out noise and prioritize high-quality features in the latent space. For imbalance class weight input, we adaptively select neighbors by class-aware scaling, mitigating class imbalance in different scenarios.
Momentum contrastive learning. To further improve the quality of pseudo-label generation and improve the discriminability of class-wise features, we introduce a contrastive learning memory bank. This memory bank ensures that the model learns more structural and high-confidence representations by storing high-quality class-specific anchor features. The memory bank maintains prototypes per class through momentum updating, with updating rate and class-wise updating embeddings . Further, the contrastive learning objective enforces structural feature discrimination.
As shown in Fig. 4, the contrastive learning module is used to pull the embedding inputs of the same classes closer to the stored anchor embeddings while pushing away from others, encouraging learning from high-quality and more discriminative features. In the update stage, the memory bank is gradually updated with the high-quality fused features. Note that the student classifier is used to predict pseudo-labels for the mapping features and guide high-quality sampling during the update stage of the memory bank. In the contrastive learning stage, the mapping feature is guided to align with high-quality stored anchor features and combined with the supervised segmentation loss during adaptation.
@@ IV-A Implementation Details
We use SynLiDAR as the synthetic dataset for synthetic-to-real adaptation and SemanticPOSS as the real-world dataset. SynLiDAR is a large-scale synthetic dataset comprising 198,396 annotated point clouds with 32 semantic classes, while SemanticPOSS consists of 2,988 annotated real-world point clouds and 14 semantic labels. For real-to-real domain adaptation, we employ SemanticKITTI, a widely used large-scale dataset that provides point-wise semantic annotations of the KITTI dataset [36]. In addition, we use the SemanticSTF dataset, which contains 2,076 manually annotated scans captured by the Velodyne HiDL64S3D LiDAR sensor from STF [37] dataset. SemanticSTF includes various adverse weather conditions, such as snow, dense fog, light fog, and rain. Following CosMix [19], we adopt the same training and validation split for , while for , we use the original validation split.